OpenStudy (anonymous):

ALGEBRA 2, PLEASE HELP ME Solve the system using elimination 2x + 3y = 5 8y = x + 26 Show steps please, I dont understand!

OpenStudy (anonymous):

2x+3y=5-------I x-8y=-26-------II then multiply the second equation by (-2) 2x+3y=5-------I -2x+16y=52-------II -------------------------- now find the sum of these two equations with each other 0x+19y=57====== y=57/19 therefore y=3 and now substitiue the [y] value the either of equations and and find x X-(8x3)=-26 -------- finally X= -26+24= X=-2